A body having a mass of 5 kg starts from rest and attains a speed of 4mpersec in 6 seconds find the force required to move the body wwith uniform acceleration in ahorizontal surface take coefficent of friction as 0.25

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From Newton’s First law of eqn. we get,
v = u +at
4 = 0 + a*6
a = 2/3 m/s^2
Force required = m*a => 5*2/3 = 10/3 N => 3.33 N
Friction force = 0.25*50 => 25/2 = 12.5 N
Net force => 12.5 – 3.33 => 9.22 approx.